A monatomic ideal gas occupies a volume Vi, which is then decreased to Vf = 1/2 Vi via an adiabatic process. Which relationship is correct for the pressures of this gas?

a) pf = pi
b) pf = 25/3 pi
c) pf = 0.55/3 pi
d) pf = 0.5 pi
Answer #1

for monoatomic ideal gas relation between Pand Vis PV\gamma=constant

Pi/Pf=(Vf/Vi)\gamma

here Vf=1/2 Vi,\gamma=5/3=1.66,then

Pi/Pf=(1/2)1.66=0.316

Pi=0.316 Pf

Pf=3.16 Pi